# Condition Weaving Meets Expert Modulation: Towards Universal and Controllable Image Generation

### v1

@@ 4.2 Evaluation Metrics
We adopt several widely used evaluation metrics in related work: FID [30], SSIM [31], CLIP-I [32, 24], CLIP-T [32, 24], and DINO [33]. FID evaluates the realism and distributional similarity between generated and real images at the feature level by extracting features using Inception V3 [34]. SSIM measures the similarity between two images in terms of luminance, contrast, and structural information. CLIP-I and CLIP-T leverage CLIP to extract visual and semantic representations. The quality of the generated images is assessed by computing feature similarity between generated and real images (CLIP-I), as well as between generated images and text prompts (CLIP-T). DINO also extracts visual features and measures the semantic structural similarity between generated and real images
4.3 Implement Details
We adopt Stable Diffusion 3.5 Medium [4]  [^1] https://huggingface.co/stabilityai/stable-diffusion-3.5-medium as the base model and initialize it with publicly available pretrained weights. The training framework is implemented using PyTorch Lightning and conducted on 8 A100 GPUs with the standard Distributed Data Parallel (DDP) strategy. No compression or acceleration techniques such as DeepSpeed [35] or FSDP [36] are employed.
During training, we use the AdamW [37] optimizer with a learning rate of 0.0001, a constant learning rate scheduler with 500 warm-up steps, and a batch size of 16. This work aims to enhance the model’s adaptability to various conditioning types. To this end, we optimize the distributed sampler to prevent batches from containing samples of the same conditioning type, which could bias the model’s optimization direction. Similar to other reproduced methods, we randomly sample 160,000 condition-description pairs for each conditioning type from the training set.
@@ 4.3 Implement Details
UniControl [10] 0.33 0.42 0.24 0.33 0.19 0.19 0.19 0.19
UniControl ++ [16] 0.39 0.41 - 0.4 0.28 0.28 - 0.28
OminiControl [13] 0.49 0.58 0.33 0.47 0.34 0.43 0.25 0.34
PixelPonder [11] 0.35 0.39 0.33 0.36 0.28 0.33 0.25 0.27
SSIM $\uparrow$ UniGen(ours) 0.5 0.55 0.39 0.48 0.34 0.43 0.28 0.35
UniControl [10] 23.34 18.23 36.69 26.09 57.63 57.31 57.84 57.59
UniControl ++ [16] 15.99 23.26 - 19.63 22.24 37.56 - 29.9
OminiControl [13] 12.35 9.38 35.03 18.92 22.92 15.34 52.0 30.09
PixelPonder [11] 24.11 21.86 30.98 25.65 24.83 20.92 65.94 37.23
FID $\downarrow$ UniGen(ours) 11.26 9.52 15.68 12.15 13.45 11.83 15.43 13.57
UniControl [10] 84.95 87.99 77.80 83.58 64.49 64.53 64.47 64.50
UniControl ++ [16] 87.04 86.75 - 86.90 73.86 72.83 - 73.35
OminiControl [13] 90.59 93.13 79.84 87.85 75.05 82.73 64.9 74.23
PixelPonder [11] 79.98 83.35 76.89 80.07 74.17 79.50 67.64 73.77
CLIP-I $\uparrow$ UniGen(ours) 89.14 91.61 82.72 87.82 76.77 84.06 71.54 77.46
UniControl [10] 18.49 18.24 19.18 18.64 19.23 19.54 19.38 19.38
UniControl ++ [16] 18.83 19.36 - 19.10 19.33 19.95 - 19.64
OminiControl [13] 19.96 20.58 20.59 20.38 19.16 19.53 19.91 19.53
PixelPonder [11] 21.56 21.75 21.4 21.57 20.62 20.63 20.19 20.48
CLIP-T $\uparrow$ UniGen(ours) 19.51 20.21 20.55 20.09 19.56 20.06 20.19 19.94
UniControl [10] 91.23 92.35 82.28 88.62 75.72 75.75 75.69 75.72
UniControl ++ [16] 93.02 91.0 - 92.0 87.99 85.96 - 86.98
OminiControl [13] 95.42 96.14 85.94 92.5 88.39 92.1 79.14 86.54
PixelPonder [11] 86.03 88.97 82.79 85.93 84.78 88.48 76.22 83.16
DINO $\uparrow$ UniGen(ours) 94.31 95.20 88.11 92.54 88.68 92.19 83.74 88.20
Table 1: We conduct a comprehensive performance comparison under Depth, Canny, and OpenPose conditions using the Subjects-200K and MultiGen-20M datasets. $\dagger$ denotes the use of FLUX [5] as the base model.
@@
Table 2: Comparison with state-of-the-art image-to-image conditional control models on the MultiGen-20M test set. $\dagger$ denotes the use of FLUX [5] as the base model.
4.4 Comparison with state-of-the-art Methods
We train and evaluate our method on Subjects-200K [13] and MultiGen-20M [16]. For fairness, we remove all test samples and randomly sample 160K instruction–target–condition image triplets per condition from the remaining data for training. We compare against traditional baselines like UniControl [10] and ControlNet++ [16], as well as recent state-of-the-art methods such as OminiControl [13] and PixelPonder [11]. Unlike OminiControl, which is built on FLUX [5], our method is based on Stable Diffusion [4].
First, as shown in Table 1, we conduct comparative evaluations on the Subjects-200K and MultiGen-20M datasets under three conditioning types: Depth, Canny, and OpenPose. It is evident that our method achieves significant improvements over existing approaches across all metrics, including SSIM, FID, CLIP-I, and DINO. Moreover, the average performance across different conditions for the same metric consistently reaches the best results. This demonstrates that our approach not only enables effective and high-quality controllable image generation but also exhibits strong generalization capabilities, achieving significant gains across various input conditions. Second, for the nine additional conditions in the MultiGen-20M dataset, as shown in Table 2, our method achieves top or near-top performance across multiple evaluation metrics. These results indicate that our approach not only ensures accurate conditional control and high image quality but also maintains strong generalization across different condition types. At the same time, we observe that our method performs relatively weaker on the CLIP-T metric. This is primarily because the proposed WeaveNet architecture reduces the reliance on prompts during conditional control. Consequently, compared to approaches that use independent branches for prompt-based global generation, our method may be less efficient at aligning image content with textual instructions, leading to suboptimal CLIP-T scores. Finally, we analyze model parameters and inference time. As shown in Table 3, traditional ControlNet [8] architectures introduce significant parameter redundancy and increased inference time as the number of condition types grows. Although OmniControl trains separate LoRA modules on top of FLUX for each condition type and avoids drastic parameter increases, its total parameter count is still several times larger than ours. This helps explain its relatively strong performance under certain conditions or metrics. Overall, our method not only maintains a compact parameter size but also achieves lower inference overhead.
@@ 4.4 Comparison with state-of-the-art Methods
Table 3: Complexity analysis between ControlNet, OminiControl and UniGen.
4.5 Ablation Studies
To comprehensively validate the effectiveness of each module, we conducted an extensive ablation study from five perspectives: the impact of the WeaveNet architecture, the roles of the Modulated Expert and Rotary Position Embedding (RoPE), the necessity of the Shared Expert, and the influence of the number of experts and conditional control layers. First, as shown in Table 4, the WeaveNet architecture significantly outperforms the original ControlNet across all three conditions in terms of SSIM, FID, CLIP-I, and DINO metrics. We also observe a slight decline in the CLIP-T metric, which is primarily due to the WeaveNet architecture potentially weakening the influence of text on the image, as it balances the joint constraints of textual and conditional visual information. Second, regarding the effectiveness of the Modulated Expert, Table 5 shows that the first row represents the conventional MoE [17, 18] design, which uses a multilayer perceptron (MLP) to process and map features. Compared to the configuration combining Modulated Expert and RoPE, there is a significant performance gap in pixel-level image generation. This indicates that using a standard MLP to handle sparsely distributed conditional visual information fails to provide precise control signals, leading to a noticeable decline in overall image reconstruction quality. Furthermore, the introduction of RoPE brings a substantial performance improvement by preserving relative spatial position information, which is often lost when processing token-level visual representations in MoE-based architectures. This effectively strengthens local conditional control. Third, regarding the Shared Expert module, as shown in Table 6, introducing the shared expert module improves FID by 1.05. Fourth, concerning the number of Modulated Experts, Table 7 shows that performance is optimal when the number of conditional experts reaches six, enabling the generation of higher-quality images. Finally, regarding the number of control layers, as illustrated in Fig. 4, both excessively deep and shallow configurations lead to significant performance degradation. Optimal generation quality is achieved when the number of control layers is set to 12.

### v3

@@ IV-B Evaluation Metrics
We adopt several widely used evaluation metrics in related work: FID [FID], SSIM [SSIM], CLIP-I [clipscore, clip], CLIP-T [clipscore, clip], and DINO [DINO]. FID evaluates the realism and distributional similarity between generated and real images at the feature level by extracting features using Inception V3 [inception]. SSIM measures the similarity between two images in terms of luminance, contrast, and structural information. CLIP-I and CLIP-T leverage CLIP to extract visual and semantic representations. The quality of the generated images is assessed by computing feature similarity between generated and real images (CLIP-I), as well as between generated images and text prompts (CLIP-T). DINO also extracts visual features and measures the semantic structural similarity between generated and real images.
To comprehensively evaluate the effectiveness of our method, we additionally incorporate the metrics from ICE-Bench[ice_bench], covering multiple dimensions including aesthetics, imaging quality, instruction-following ability, reference consistency, and controllability. From the aesthetics perspective, we adopt AES (Aesthetic Score), which uses an aesthetic predictor to assess the overall visual appeal of generated images. This metric considers composition, color usage, consistency, and naturalness, providing an effective measure of generative quality and visual attractiveness. For imaging quality, we include IMG (Imaging Quality), which evaluates whether an image is sharp and free of distortions or exposure issues. To assess instruction adherence, we use CLIP-Cap, which computes the similarity between images and semantic descriptions in the feature space, offering a direct measure of the model’s ability to interpret and follow textual prompts. Finally, we assess pixel-level differences using L1-based metrics. L1 Raw measures the mean absolute pixel difference between the generated and ground-truth images, reflecting pixel-level alignment. L1 Color evaluates color restoration and color-shift magnitude, indicating the color discrepancy relative to the ground truth. L1 Cond computes the pixel-level difference between the condition image extracted from the generated output and the reference condition image. It reflects how well the generated image spatially fits the reference condition. Following the evaluation protocol of ICE-Bench, all reported values for L1 Color and L1 Cond are presented as $1-L1_{Color}$ and $1-L1_{Cond}$, respectively.
To further align with standard evaluation practices in image generation tasks, we extend our assessment using additional image quality metrics. Following the methodology adopted in IQA-Pytorch [^6] https://github.com/chaofengc/IQA-PyTorch, we evaluate performance across MUSIQ [musiq], MAN-IQA [maniqa], CLIP-IQA [clip-iqa], LPIPS [lpips], LAION-Aesthetics [laion_aes], PSNR, NIMA [nima], TOPIQ-NR [topiq], and CLIPScore [clip]. Among these, MUSIQ, CLIP-IQA, LAION-Aesthetics, and CLIPScore share similar purposes with several ICE-Bench metrics, characterizing aesthetic quality, imaging fidelity, and semantic or visual alignment. In addition, we introduce LPIPS, PSNR, NIMA, and TOPIQ-NR, which evaluate perceptual distance, pixel-level error, human aesthetic preference, and no-reference image quality, respectively, enabling a more comprehensive assessment of the generated images.
IV-C Experimental Setting
To comprehensively evaluate the effectiveness of our method, we conduct experiments from two perspectives: single-condition control and multi-condition joint control. Since the MultiGen-20M dataset contains 12 types of independent conditioning signals, we use this dataset only for single-condition image generation and compare our method with conventional approaches. The Subjects-200k dataset provides four types of conditioning signals, among which three can be used independently. Therefore, we use the depth, canny, and openpose conditions from Subjects-200k to train and evaluate the single-condition control task. To further assess our method, we follow the experimental design of UniCombine [Unicombine] and construct multi-condition control tasks by combining: Depth + Canny, Subject + Depth, and Subject + Canny. These combinations are used to train and evaluate the joint-condition control setting. In Section IV-E1, we evaluate and analyze the generation results for single-condition control tasks on the MultiGen-20M and Subjects-200k datasets. In Section IV-E2, we focus on evaluating and analyzing the generation results for multi-condition control tasks on the Subjects-200k dataset.
@@ IV-C Experimental Setting
UniControl [Unicontrol] 0.33 0.42 0.24 0.33 0.19 0.19 0.19 0.19
Controlnet ++ [cn++] 0.39 0.41 - 0.4 0.28 0.28 - 0.28
OminiControl [ominicn] 0.49 0.58 0.33 0.47 0.34 0.43 0.25 0.34
PixelPonder [Pixelponder] 0.35 0.39 0.33 0.36 0.28 0.33 0.25 0.27
UniGen(ours) 0.5 0.55 0.39 0.48 0.34 0.43 0.28 0.35
SSIM $\uparrow$ UniGen(ours) 0.5 0.59 0.37 0.49 - - - -
UniControl [Unicontrol] 23.34 18.23 36.69 26.09 57.63 57.31 57.84 57.59
Controlnet++ [cn++] 15.99 23.26 - 19.63 22.24 37.56 - 29.9
OminiControl [ominicn] 12.35 9.38 35.03 18.92 22.92 15.34 52.0 30.09
PixelPonder [Pixelponder] 24.11 21.86 30.98 25.65 24.83 20.92 65.94 37.23
UniGen(ours) 11.26 9.52 15.68 12.15 13.45 11.83 15.43 13.57
FID $\downarrow$ UniGen(ours) 11.51 8.32 26.51 15.45 - - - -
UniControl [Unicontrol] 84.95 87.99 77.80 83.58 64.49 64.53 64.47 64.50
Controlnet++ [cn++] 87.04 86.75 - 86.90 73.86 72.83 - 73.35
OminiControl [ominicn] 90.59 93.13 79.84 87.85 75.05 82.73 64.9 74.23
PixelPonder [Pixelponder] 79.98 83.35 76.89 80.07 74.17 79.50 67.64 73.77
UniGen(ours) 89.14 91.61 82.72 87.82 76.77 84.06 71.54 77.46
CLIP-I $\uparrow$ UniGen(ours) 90.94 93.67 82.47 89.03 - - - -
UniControl [Unicontrol] 18.49 18.24 19.18 18.64 19.23 19.54 19.38 19.38
Controlnet++ [cn++] 18.83 19.36 - 19.10 19.33 19.95 - 19.64
OminiControl [ominicn] 19.96 20.58 20.59 20.38 19.16 19.53 19.91 19.53
PixelPonder [Pixelponder] 21.56 21.75 21.4 21.57 20.62 20.63 20.19 20.48
UniGen(ours) 19.51 20.21 20.55 20.09 19.56 20.06 20.19 19.94
CLIP-T $\uparrow$ UniGen(ours) 22.28 22.38 22.21 22.29 - - - -
UniControl [Unicontrol] 91.23 92.35 82.28 88.62 75.72 75.75 75.69 75.72
Controlnet++ [cn++] 93.02 91.0 - 92.0 87.99 85.96 - 86.98
OminiControl [ominicn] 95.42 96.14 85.94 92.5 88.39 92.1 79.14 86.54
PixelPonder [Pixelponder] 86.03 88.97 82.79 85.93 84.78 88.48 76.22 83.16
UniGen(ours) 94.31 95.20 88.11 92.54 88.68 92.19 83.74 88.20
DINO $\uparrow$ UniGen(ours) 95.37 96.44 88.14 93.32 - - - -
TABLE I: We conduct a comprehensive performance comparison under Depth, Canny, and OpenPose conditions using the Subjects-200K and MultiGen-20M datasets. $\dagger$ denotes the use of FLUX [flux.1] as the base model.
IV-D Implement Details
We adopt Stable Diffusion [sd35] [^7] https://huggingface.co/stabilityai/stable-diffusion-3.5-medium and Flux .1888https://huggingface.co/black-forest-labs/FLUX.1-schnell [flux.1] as the base model and initialize it with publicly available pretrained weights. The training framework is implemented using PyTorch Lightning and conducted on 8 A100 GPUs with the standard Distributed Data Parallel (DDP) strategy. When training models based on Flux .1, the parameter count is significantly larger than that of Stable Diffusion. Therefore, we adopt the DeepSpeed ZeRO-3 [deepspeed] configuration provided by Accelerate  [^9] https://github.com/huggingface/accelerate to partition and offload parameters, which helps reduce the risk of GPU memory overflow.
During training, we use the AdamW [adamw] optimizer with a learning rate of 0.0001, a constant learning rate scheduler with 500 warm-up steps, and a batch size of 16. This work aims to enhance the model’s adaptability to various conditioning types. To this end, we optimize the distributed sampler to prevent batches from containing samples of the same conditioning type, which could bias the model’s optimization direction. Similar to other reproduced methods, we randomly sample 160,000 condition-description pairs for each conditioning type from the training set.
@@ IV-D Implement Details
TABLE II: Comparison with state-of-the-art image-to-image conditional control models on the MultiGen-20M test set. $\dagger$ denotes the use of FLUX [flux.1] as the base model.
Task Method AES $\uparrow$ IMG $\uparrow$ CLIP Cap $\uparrow$ L1 Raw $\downarrow$ L1 Color $\uparrow$ L1 Cond $\uparrow$ Final Score $\uparrow$
Depth ControlNet 5.34 51.52 0.26 0.16 0.83 0.93 0.59
UniControl 5.19 48.82 0.27 0.24 0.76 0.92 0.58
Controlnet ++ 5.30 50.48 0.26 0.20 0.80 0.93 0.59
OminiControl 5.42 51.51 0.25 0.15 0.86 0.93 0.59
PixelPonder 4.96 48.71 0.24 0.22 0.80 0.88 0.56
UniGen(ours) 5.34 50.11 0.26 0.15 0.85 0.93 0.59
UniGen(ours) 5.48 52.78 0.25 0.13 0.87 0.93 0.64
Canny ControlNet 5.36 50.16 0.26 0.16 0.84 0.93 0.59
UniControl 5.35 47.78 0.26 0.22 0.79 0.92 0.58
Controlnet ++ 5.43 42.71 0.26 0.20 0.80 0.91 0.57
OminiControl 5.43 49.77 0.25 0.13 0.88 0.94 0.59
PixelPonder 4.85 49.25 0.24 0.22 0.81 0.90 0.56
UniGen(ours) 5.31 49.74 0.26 0.14 0.87 0.94 0.59
UniGen(ours) 5.45 51.34 0.25 0.11 0.90 0.93 0.64
Openpose ControlNet 5.37 49.93 0.26 0.21 0.78 0.97 0.59
UniControl 5.12 47.42 0.27 0.29 0.71 0.97 0.56
Controlnet ++ - - - - - - -
OminiControl 5.39 48.11 0.26 0.21 0.79 0.98 0.58
PixelPonder 5.00 47.60 0.24 0.23 0.80 0.98 0.56
UniGen(ours) 5.39 48.28 0.27 0.22 0.79 0.98 0.58
UniGen(ours) 5.49 51.71 0.26 0.20 0.81 0.98 0.64
TABLE III: Performance evaluation of generation quality under three conditioning settings on the Subjects-200K dataset using the ICE-Bench evaluation protocol. $\dagger$ denotes the use of FLUX [flux.1] as the base model.
Task Method MUSIQ $\uparrow$ MAN-IQA $\uparrow$ CLIP-IQA $\uparrow$ LPIPS $\downarrow$ LAION-AES $\uparrow$ PSNR $\uparrow$ NIMA $\uparrow$ TOPIQ-NR $\uparrow$ CLIPSCORE $\uparrow$
Depth ControlNet 58.96 0.49 0.56 0.55 5.62 12.75 5.00 0.70 0.79
UniControl 59.98 0.48 0.52 0.63 5.74 10.28 4.87 0.71 0.81
Controlnet ++ 61.00 0.49 0.54 0.58 5.74 11.59 4.99 0.71 0.80
OminiControl 60.21 0.49 0.56 0.50 5.62 14.16 4.89 0.71 0.77
PixelPonder 58.91 0.46 0.51 0.66 5.71 10.89 4.98 0.69 0.74
UniGen(ours) 58.78 0.49 0.54 0.52 5.69 13.45 4.97 0.71 0.79
UniGen(ours) 60.66 0.52 0.58 0.50 5.75 14.57 5.04 0.73 0.78
Canny ControlNet 58.16 0.48 0.55 0.51 5.63 13.32 4.98 0.70 0.79
UniControl 57.21 0.45 0.49 0.58 5.68 11.06 4.82 0.69 0.80
Controlnet ++ 53.38 0.39 0.45 0.60 5.66 11.43 4.81 0.68 0.80
OminiControl 59.22 0.48 0.55 0.45 5.60 15.30 4.90 0.70 0.76
PixelPonder 60.39 0.48 0.52 0.61 5.63 11.13 5.00 0.70 0.74
UniGen(ours) 58.76 0.48 0.54 0.48 5.67 14.37 4.99 0.72 0.79
UniGen(ours) 59.89 0.52 0.58 0.42 5.71 16.19 5.03 0.73 0.78
Openpose ControlNet 57.49 0.47 0.55 0.66 5.58 10.75 5.00 0.69 0.80
UniControl 59.38 0.45 0.51 0.75 5.76 8.56 4.76 0.71 0.82
Controlnet ++ - - - - - - - - -
OminiControl 57.58 0.44 0.53 0.68 5.57 11.20 4.72 0.68 0.79
PixelPonder 57.62 0.45 0.50 0.69 5.70 10.41 4.95 0.68 0.75
UniGen(ours) 56.78 0.47 0.54 0.65 5.70 10.86 4.95 0.70 0.81
UniGen(ours) 59.83 0.52 0.57 0.64 5.84 11.69 4.92 0.72 0.79
TABLE IV: Performance evaluation of generation quality under three conditioning settings on the Subjects-200K dataset using the IQA-Pytorch assessment framework. $\dagger$ denotes the use of FLUX [flux.1] as the base model.
@@
TABLE V: Complexity analysis between ControlNet, UniControl, OminiControl and UniGen. $\dagger$ denotes the use of FLUX [flux.1] as the base model.
Evaluate on traditional metrics
Task Method Base Model PNSR $\uparrow$ SSIM $\uparrow$ FID $\downarrow$ CLIP-I $\uparrow$ CLIP-T $\uparrow$ LPIPS $\downarrow$ DINO $\uparrow$ Param (B) $\downarrow$
Depth + Canny UniCombine [Unicombine] Flux .1 [flux.1] 13.66 0.45 11.09 92.02 29.90 0.56 97.03 11.93
UniGen(Ours) SD 3.5 [sd35] 15.65 0.62 9.03 92.73 30.45 0.43 96.54 3.41
Subject + Depth UniCombine [Unicombine] Flux .1 [flux.1] 13.18 0.42 11.56 91.25 29.89 0.58 96.27 11.93
UniGen(Ours) SD 3.5 [sd35] 13.28 0.48 11.69 89.36 30.72 0.52 94.28 3.41
Subject + Canny UniCombine [Unicombine] Flux .1 [flux.1] 13.66 0.45 9.51 92.52 29.71 0.58 96.63 11.93
UniGen(Ours) SD 3.5 [sd35] 14.59 0.57 9.66 91.86 30.58 0.47 95.27 3.41
Evaluate on ICE Bench [ice_bench] (ICCV 2025)
Task Method Base Model AES $\uparrow$ IMG $\uparrow$ CLIP Cap $\uparrow$ CLIP Src $\uparrow$ L1 Raw $\downarrow$ L1 Color $\uparrow$ Final Score $\uparrow$ Param (B) $\downarrow$
Depth + Canny UniCombine [Unicombine] Flux .1 [flux.1] 5.37 58.16 0.25 - 0.14 0.86 0.60 11.93
UniGen(Ours) SD 3.5 [sd35] 5.46 48.84 0.26 - 0.12 0.88 0.59 3.41
Subject + Depth UniCombine [Unicombine] Flux .1 [flux.1] 5.04 52.76 0.25 0.72 0.15 0.84 0.57 11.93
UniGen(Ours) SD 3.5 [sd35] 5.32 48.09 0.26 0.68 0.16 0.84 0.57 3.41
Subject + Canny UniCombine [Unicombine] Flux .1 [flux.1] 5.09 55.52 0.25 0.72 0.14 0.85 0.58 11.93
UniGen(Ours) SD 3.5 [sd35] 5.36 48.76 0.26 0.70 0.13 0.87 0.58 3.41
TABLE VI: Performance evaluation of multi-condition generation results using both traditional metrics and the ICE-Bench evaluation protocol.
IV-E Comparison and Analysis with state-of-the-art Methods
IV-E1 Comparison and Analysis of Single-Condition Controllable Image Generation Performance
We train and evaluate our method on Subjects-200K [ominicn] and MultiGen-20M [cn++]. For fairness, we remove all test samples and randomly sample 160K instruction–target–condition image triplets per condition from the remaining data for training. We compare against traditional baselines like UniControl [Unicontrol] and ControlNet++ [cn++], as well as recent state-of-the-art methods such as OminiControl [ominicn] and PixelPonder [Pixelponder].
First, as shown in Table I, we conducted a comparative evaluation on the Subjects-200K and MultiGen-20M datasets for three types of conditions (Depth, Canny, and OpenPose) using traditional metrics. It is clear and intuitive that our method outperforms existing approaches across all three condition types. Specifically, when combined with the current best base model, Flux .1 [flux.1], our method demonstrates a significant advantage, achieving the highest average performance across all metrics and substantially surpassing other methods based on the same base model. Meanwhile, when applied to the smaller Stable Diffusion 3.5 model, our method still achieves competitive, near-optimal performance. This indicates that our approach not only exhibits strong generalization capability but also effectively enables controllable image generation according to different condition types.
Next, as shown in Table II, we compare image generation quality across the remaining nine condition types on the MultiGen-20M dataset. Even using the smaller Stable Diffusion as the base model, our method achieves the best or second-best performance in multiple evaluation metrics across these nine condition types, in some cases significantly outperforming methods based on Flux .1. Overall, our method ensures precise condition control and high-quality image content while maintaining strong generalization across diverse condition types.
Considering that traditional metrics may be imperfect in evaluating image quality and aesthetics, we further conducted comprehensive performance assessments under the ICE-Bench [ice_bench] and IQA-Torch frameworks. As shown in Table III, ICE-Bench provides multi-faceted evaluation, including aesthetics, image quality, instruction adherence, and condition control. Table III clearly shows that our method achieves the best or second-best performance under two different base models. Notably, in terms of aesthetics and image quality, our method significantly outperforms the next-best approach. For example, under Depth conditions, compared to OminiControl with the same base model, our method improves image quality by +1.27; under Canny and OpenPose conditions, improvements of +1.5 and +3.6 are observed, respectively.
Similarly, as shown in Table IV, evaluations conducted with IQA-PyTorch-based image quality metrics indicate that our method achieves the best performance across multiple dimensions, including aesthetics, image quality, and perceptual distance, substantially surpassing related methods based on similar base models. However, in prior evaluations, we noticed that our method is relatively less competitive in metrics related to adherence to textual instructions, generally ranking second rather than first. We attribute this to the iterative interaction mechanism in our WeaveNet architecture for condition feature fusion and injection. Compared to traditional ControlNet [controlnet] or LoRA-based [lora, ominicn] methods, WeaveNet focuses more on processing visual condition information during intermediate steps, which reduces the influence of the original textual instructions and decreases reliance on prompts during condition control. Therefore, compared with methods that employ independent branches for prompt-based global generation, our method may perform slightly lower in aligning image content with text instructions, leading to suboptimal CLIP-T and CLIPScore results. This issue is further discussed in Section VI.
Finally, we analyze model parameters and inference time. As shown in Table V, the traditional ControlNet [controlnet] architecture introduces significant parameter redundancy, and inference time increases with the number of condition types. This is primarily because ControlNet relies on parallel computation with relatively independent processing branches; after independently extracting condition features via ControlNet branches, the features are interactively injected into the base model, which inevitably increases inference time. OminiControl, based on the Flux model, trains separate LoRA modules [lora] for each condition type. While this introduces fewer additional parameters, aligning with LoRA’s design goals, it is limited in feature processing and overall performance. In contrast, our method, even when applied to a smaller base model with fewer parameters and lower intrinsic capability than Flux, achieves superior performance across multiple evaluation metrics and condition types.
Although our design introduces more parameters compared to LoRA [lora], we only require a single independent ControlNet architecture with additional expert control units as the number of conditions increases. This keeps the parameter growth minimal while maintaining comparable parameter increments to LoRA. Importantly, our approach achieves significantly greater performance gains compared to LoRA-based methods. As shown in Table II, when scaling to 12 tasks, our model introduces an additional 0.59B parameters, while OminiControl adds 0.14B. Although our parameter increase is four times larger, our method is built on a much smaller 2.5B-parameter Stable Diffusion backbone yet achieves higher performance than OminiControl, which is based on the larger Flux backbone. For example, in terms of FID, our model outperforms OminiControl by 4.37, 7.66, 6.92, 19.32, and 50.84 on the Hed, HedSketch, Normal, Seg, Bbox, and Outpainting tasks, respectively. Note that these tasks are more complex and challenging than Depth, Canny, and OpenPose, making them a stronger indicator of the model’s generative quality and controllability.
IV-E2 Comparison and Analysis of Multi-Condition Controllable Image Generation Performance
To provide a more comprehensive evaluation of our approach, we further extend it to multi-condition joint control tasks. Compared with single-condition control, multi-condition joint control introduces more constraints and imposes stronger requirements on spatial alignment and subject consistency. Although one might expect that adding more conditions would simplify generation and improve performance, the opposite is true. Multi-condition joint control strictly restricts the model’s output—for example, enforcing consistency across subjects, spatial depth, and texture boundaries. As a result, many image generation methods that perform well under single-condition control fail to produce valid outputs in this setting.
Since this is an emerging task, there are few comparable reference methods. Therefore, we follow the evaluation setup of the UniCombine [Unicombine] and conduct assessments under two different evaluation protocols. Considering that traditional approaches lose controllability in this task and cannot produce usable results, we do not include them in our comparisons. Instead, we focus on evaluating against UniCombine, the most recent and competitive method. Note that UniCombine, similar to OminiControl, is built on the larger and more capable Flux.1 base model, and achieves multi-condition joint control by independently training multiple LoRA components.
As shown in Table VI, under the standard evaluation metrics, our method consistently achieves better performance across multiple metrics in three multi-condition control tasks. It should be noted that our model uses substantially fewer parameters than UniCombine, reducing the parameter count by 8.52B. Despite this reduction, the performance gap on several metrics remains minimal. Even in cases where our method is slightly weaker, the difference is within a margin of 1%.
We further conduct a comprehensive evaluation using ICE-Bench. The results show that our approach achieves advantages in several aspects, including aesthetics and conditional controllability. Notably, our method demonstrates a significant improvement in aesthetic scores across all three multi-condition control tasks compared to UniCombine. As discussed in Section IV-G2, multi-condition tasks impose stronger constraints, making our model produce more natural outputs. In contrast, UniCombine tends to overfit the conditional information, which leads to reduced aesthetic quality.
In general, we conducted comprehensive evaluations across multiple datasets, tasks, and metric systems. Our method consistently achieves significant improvements and, in many cases, even exceeds approaches with larger parameter counts and stronger base-model capabilities. This demonstrates that our method can effectively follow conditioning instructions while generating content that better aligns with aesthetic evaluation standards. Moreover, when extending from a single condition to multiple conditions, our approach introduces only a small increase in parameters—comparable to LoRA-based methods—yet delivers substantially greater performance gains.
IV-F Ablation Studies
To comprehensively validate the effectiveness of each module, we conducted an extensive ablation study from five perspectives: the impact of the WeaveNet architecture, the roles of the Modulated Expert and Rotary Position Embedding (RoPE), the necessity of the Shared Expert, and the influence of the number of experts and conditional control layers.
IV-F1 Ablation Analysis of the WeaveNet Architecture
In traditional ControlNet, conditional feature processing is performed in parallel with the backbone network. Specifically, an independent copy of the backbone is used as a conditional control unit to extract features, which are then added to the features of each corresponding layer of the backbone to achieve global conditional feature injection. First, we evaluate using only the ControlNet architecture, without incorporating our proposed CoMoE module. As shown in the upper part of Table VII, without CoMoE, simply maintaining the parallel design of ControlNet yields inferior performance compared to our method. For example, in terms of FID, our method outperforms the traditional ControlNet across Depth, Canny, and Openpose by 1.34, 0.82, and 0.68, respectively.
To further validate the advantage of the WeaveNet architecture, we also modified WeaveNet to use the traditional ControlNet-style feature injection, keeping the backbone network and conditional feature network independent. As shown in the lower part of Table VII, even under the WeaveNet framework, switching to the parallel feature injection of ControlNet still produces a significant performance gain. This demonstrates that, compared to traditional ControlNet, WeaveNet better integrates the backbone and conditional control networks, enabling more precise and effective conditional control. In contrast, traditional ControlNet tends to process the two branches independently, which limits both feature fusion and accurate conditional control.
Wo. CoMoE